OpenStudy (anonymous):

evaluate 36^3/2 please help!!!

OpenStudy (anonymous):

your exponent has a denominator of 2 and a numerator of 3 the denominator means the root what is the square root of 36 ?

OpenStudy (anonymous):

6? what would be the answer

OpenStudy (anonymous):

no that is a step on the say now you have to cube 6, i.e. compute \(6^3=6\times 6\times 6\) then you will be done

OpenStudy (anonymous):

step on the WAY